Use a triple integral to find the volume of the given solid. The solid enclosed by the cylinder y = x^2 and the planes z = 0 and y + z = 1.


1
Expert's answer
2021-07-14T14:40:09-0400

y = x2

z = 0 and y + z = 1


The limits of x, y and z are

x = -1 to 1

y = x2 to 1

z = 0 to 1 - y


Let x2 = k



The volume of the cylinder is given by


V = "\int"-1"\int" k"\int" 01-y dz dy dx



V = "\int"-1"\int" k1  ( 1 - y ) dy dx


V = "\int"-1( y - "\dfrac{y^2}{2}" )k 1 dx


V = "\int"-1( 1 - k - "\dfrac{1}{2}""\dfrac{k^2}{2}" ) dx





On substituting the value of k, we have


V = "\int"-1( 1 - x2 - "\dfrac{1}{2}" + "\dfrac{x^4}{2}" ) dx


V = "\int"-1"\dfrac{1}{2}" - x"\dfrac{x^4}{2}" ) dx


V = ( "\dfrac{1}{2}"x - "\dfrac{x^3}{3}" + "\dfrac{x^5}{10}" )-1 1


V = [ "\dfrac{1}{2}"(1 + 1) - "\dfrac{(1+1)^3}{3}" + "\dfrac{(1+1)^5}{10}" ]


V = [ 1 - "\dfrac{8}{3}" + "\dfrac{32}{10}" ]


V = 1.533 cubic units.